### v2

# Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?

Abstract
Uncovering hidden symbolic laws from time series data, as an aspiration dating back to Kepler’s discovery of planetary motion, remains a core challenge in scientific discovery and artificial intelligence. While Large Language Models show promise in structured reasoning tasks, their ability to infer interpretable, context-aligned symbolic structures from time series data is still underexplored. To systematically evaluate this capability, we introduce SymbolBench, a comprehensive benchmark designed to assess symbolic reasoning over real-world time series across three tasks: multivariate symbolic regression, Boolean network inference, and causal discovery. Unlike prior efforts limited to simple algebraic equations, SymbolBench spans a diverse set of symbolic forms with varying complexity. We further propose a unified framework that integrates LLMs with genetic programming to form a closed-loop symbolic reasoning system, where LLMs act both as predictors and evaluators. Our empirical results reveal key strengths and limitations of current models, highlighting the importance of combining domain knowledge, context alignment, and reasoning structure to improve LLMs in automated scientific discovery.
Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?
Zewen Liu1, Juntong Ni1, Xianfeng Tang2, Max S.Y. Lau1, Wenpeng Yin3, Wei Jin1 1Emory University, 2Amazon, 3Penn State University {zewen.liu, juntong.ni, msy.lau, wei.jin}@emory.edu, xianfengtang@outlook.com, wenpeng@psu.edu
1 Introduction
Centuries ago, Johannes Kepler revolutionized our understanding of the cosmos by discovering the laws of planetary motion Gentner (2002). Through meticulous analysis and rigorous reasoning over astronomical observations, captured as time series data of planetary positions, Kepler derived precise mathematical relationships that explained the complex, time-dependent dynamics of celestial orbits. Automating such a challenging process, where hidden symbolic laws are extracted from time series, is a long-standing aspiration in artificial intelligence Reddy and Shojaee (2025). However, achieving this automation presents fundamental reasoning-driven challenges, as time series data encapsulates dynamic behaviors and temporal dependencies that demand abstraction, generalization, and reasoning beyond mere pattern recognition to uncover underlying symbolic structures.
Recent advances in Large Language Models (LLMs) and Multimodal LLMs (MLLMs) show strong performance in complex reasoning tasks Wang et al. (2024), but their ability to extract symbolic laws from time series remains underexplored. Traditional methods in symbolic regression, such as genetic programming Makke and Chawla (2024), often prioritize data fit at the expense of interpretability. Meanwhile, recent attempts Merler et al. (2024); Li et al. (2024); Shojaee et al. (2024) to use LLMs in this domain have been limited; they employ them merely as domain-agnostic function generators, overlooking the models’ core potential for deeper, theory-aligned reasoning. This shallow integration often results in proposed equations that lack contextual relevance. Additionally, most work focuses on algebraic expressions, neglecting other symbolic forms like logical formulas Zhang et al. (2024) and causal relations Assaad et al. (2022).
To address these gaps and provide further insights into the symbolic reasoning ability of LLMs for time series, a comprehensive benchmark is urgently required. Our work is guided by four key objectives: (a) Real-world relevance: To use real-world time series with ground-truth symbolic structures. (b) Task difficulty: To incorporate a diverse form of symbolic structures with varying complexity. (c) Scale and balance: To ensure a sufficiently large and balanced sample distribution across tasks. (d) Unified framework: To provide a unified framework to execute various tasks and establish connections with task-specific baselines.
To realize these objectives, we present three primary contributions.
(I) SymbolBench, a comprehensive benchmark designed to rigorously evaluate the symbolic reasoning capabilities of LLMs over time series with rich real-world contextual descriptions. It uniquely spans three core tasks covering major types of time series data: (a) Multivariate symbolic regression for continuous data to recover complex equations like coupled Ordinary Differential Equations (ODEs). (b) Boolean network inference Zhang et al. (2024) for discrete systems to identify logical rules. (c) Causal discovery Assaad et al. (2022) for multivariate data to uncover structured causal graphs. Each task includes challenging, real-world examples from domains such as biology, physics, and healthcare, with varying dimensionality and difficulty. To ensure both quality and coverage, we curate representative subsets from large databases, yielding a benchmark that is broader, more challenging, and more balanced than prior efforts.
(II) Unified Symbolic Reasoning Framework that enables the context-aware and knowledge-rich LLMs to play the dual role of predictors and judges in the process of hypothesis generation, testing, and refinement, while optionally including efficient genetic programming tools in a hybrid way.
(III) Critical empirical insights into the current strengths and limitations of LLMs and MLLMs in temporal symbolic reasoning. (a) LLMs surpass traditional baselines on multivariate symbolic regression and causal discovery, but fall short on Boolean network inference; (b) LLMs are able to perform a certain level of reasoning on the three tasks, with properly increasing test-time compute bringing moderate improvement. (c) Introducing context not only improves performance but also potentially guides the selection of generalizable symbolic structures. (d) Combining LLMs in complementary roles with genetic programming further boosts performance.
2 Related Work
Symbolic Expression Discovery. Symbolic regression (SR) aims to recover interpretable equations from time series data. Classical methods like Genetic Programming and sparse optimization (e.g., SINDy Brunton et al. (2016), PySR Cranmer (2024)) prioritize accuracy and parsimony but face scalability challenges. Recent deep learning models (e.g., ODEformer d’Ascoli et al. (2023), TPSR Shojaee et al. (2023)) improve efficiency by treating SR as a translation task, though they require pretraining and lack iterative refinement. Related fields like Boolean network inference and causal discovery also seek symbolic structures from time series. More details are in Appendix A.
LLM Symbolic Reasoning. Due to the strong in-context learning and reasoning ability of LLMs that allow them to adapt to various tasks, they are able to perform logical inferenceAhn et al. (2024); Wang and Chen (2023), and temporal symbolic reasoning Fang et al. (2024). Despite the strong ability of LLMs, current research for SR tasks only applies them as function generators without a reasoning process. Though the generated functions may achieve a high fitting score, the reasoning process remains unknown to us and may fail to align with the context, with little real-world meaning. In addition, the current research focus has also skewed toward algebraic equations, with logical rules or causal relations remaining underexplored.
3 SymbolBench Dataset
To rigorously evaluate the reasoning abilities of LLMs on time series related science discovery, we introduce SymbolBench, a curated benchmark that aims to uncover symbolic structures from time series. The dataset spans diverse domains (e.g., physics, biology) and is structured around three core categories of symbolic structures. More details of the dataset are provided in Appendix B.
Coupled Differential Equations (CDEs). CDEs represent dynamic systems, yielding continuous and multivariate time series data. Let $x_{i}$ denote the $i$-th variable of a multivariate time series, a coupled differential equation can be described as the following: $\frac{dx_{i}}{dt}=f_{i}(x_{1},x_{2},...,x_{n}),i=1,...,n$, where the symbolic structures $f_{i}$ describe dynamic interactions among state variables. The corresponding time series data consists of numerical solutions $\{x_{i}(t)\}$ over time, generated from initial conditions and system parameters. This setting reflects real-world dynamical systems in physics and engineering, where the complexity arises from variable interdependence. While the previous benchmark dataset, ODEbench d’Ascoli et al. (2023), contains coupled ODEs from 1 dimension to 4 dimensions, the number of samples is small, and the class of dimensions is heavily imbalanced, with only three 4-dimensional ODEs. In this study, we further enrich ODEbench with more high-dimensional ODEs and provide a balanced dataset with over 156 samples. Each sample is accompanied by the variable descriptions and the domain name if available, as shown in Appendix B.
Boolean Networks (BNs). Multivariate time series with discrete values are also seen in the scientific domains. Derived from models used in systems biology, particularly in gene regulatory and signaling networks, Boolean networks represent each variable as a binary node whose state evolves according to logical expressions: $x_{i}^{(t+1)}=f_{i}(x_{1}^{(t)},x_{2}^{(t)},...,x_{n}^{(t)}),x_{i}\in\{0,1\},$ where $f_{i}$ is a logical function composed of AND, OR, NOT, etc. The time series data is a sequence of binary vectors over discrete time steps, representing the dynamic evolution of the system. This setup emphasizes symbolic logic reasoning, state transitions, and rule discovery from temporal traces of binary states. In this study, we provide a curated subset of 65 Boolean networks from BioDivine Pastva et al. (2023). For each sample, we provide a short description of the domain and the name of each variable.
Structured Causal Models (SCMs). Beyond specifying a specific data-generating process (e.g., via mathematical functions), studying causal dependencies among time series variables is valuable for uncovering interdependencies directly from raw data, akin to the broader task of causal discovery in temporal settings. This approach models systems using Structural Causal Models (SCMs), which can be expressed as a directed graph. For each variable $x_{i}$, the goal is to identify its parent variables $x_{j}$ along with their corresponding time lags $l$, such that $x_{j}$ at time tlt-l causally influences $x_{i}$ at time $t$: $x_{i}\leftarrow\{(x_{j},l)\mid x_{j}\in X,;l\in[1,M]\},$ where $X$ denotes the set of all variables and $M$ is the maximum considered lag. As the number of variables and potential lag intervals increases, the search space for an optimal SCM grows exponentially, making discovery more challenging. In this work, we extract SCMs from the CDEs in our curated dataset, as well as from additional CDEs in the Physiome database involving more than three variables, using functional analysis. Each sample is annotated with its corresponding SCM, resulting in 190 samples.
Table 1: Comparison of SymbolBench with existing benchmarks for LLMs across symbolic structures, evaluation setups, contextual data types, and reasoning.
Feature LLM-SRBench ODEBench RealTCD Ours
Symbolic Structures
   Scientific Eq. 128 63 156
   Logical Exp. 65
   SCM 2 190
Evaluation
   ID/OOD $\checkmark$ $\checkmark$ $\checkmark$
   Reasoning $\checkmark$
Context
   Textual $\checkmark$ $\checkmark$
   Multimodal $\checkmark$
Figure 1: Iterative refinement framework. Candidate proposals are generated using either an LLM-as-Predictor or genetic programming operations. Each round of candidates undergoes quantitative and qualitative evaluation via validation tools and an LLM-as-Judge. Scored candidates are stored in a history pool, and a context manager decides contextual information for the next round.
4 SymbolBench Reasoning Framework
Given an input time series with T time points $\{\mathbf{x}^{i}\}_{i=1}^{T}$, where $\mathbf{x}^{i}\in\mathbb{R}^{D}$ and D is the number of dimensions, our framework aims to generate a subset of symbolic structures for each time series. Figure 1 sketches the closed-loop workflow that combines LLMs into the automatic pipeline. The process starts with Proposal Generation (Sec. 4.1), which generates candidate expressions alongside their reasoning path if available. Then, the raw output is cleaned and used for Verification (Sec. 4.2), which assigns scores for each candidate. Candidates are then stored in a history pool. Finally, before the next round of generation, the Context Manager (Sec. 4.3) extracts candidates from the history pool based on certain rules and provides them as the context for the next round proposal generation. The loop repeats until either the stopping criterion is met or the budget exceeds the limit.
4.1 Proposal Generation
LLM-as-Predictor. To leverage the knowledge embedded in LLMs, recent studies have explored the direct generation of equations. In this work, we frame this approach as LLM-as-Predictor. To thoroughly assess both the capabilities of LLMs and the impact of iterative refinement, external context, and chain-of-thought reasoning, we introduce four distinct prompting strategies to evaluate the performance of LLMs/MLLMs comprehensively. (a) Naive prompt instructs the model to generate mathematical expressions without providing any contextual or historical information. (b) Base prompt builds upon the Naive prompt by incorporating previously generated expressions from the history pool, filtered by the context manager. (b) Context prompt further enhances the Base prompt by adding relevant contextual information, such as variable descriptions. (c) CoT prompt extends the Context prompt by enabling the model to perform step-by-step reasoning. The reasoning process is also recorded and subjected to qualitative verification. Details of all prompts are provided in Appendix I.
Hybrid Method. While the generation of symbolic structures is central to the pipeline, this part can also be addressed using operations from traditional genetic programming. In the context of CDEs and BNs, this involves applying operations like mutation and crossover on expression trees to produce new candidates. Since LLMs may still contribute to other components of the framework, e.g., verification, we phrase such a combination as the Hybrid Method, and further discuss it in Appendix F.
4.2 Verification
Verification in symbolic structure discovery assesses three key aspects: numerical fitness, symbolic fitness, and context alignment. While numerical fitness can be quantified, symbolic fitness and context alignment are harder to measure without ground truth. To address this, we introduce a dual-verification strategy combining standard quantitative metrics with rubric-based qualitative evaluation by LLMs-as-Judges.
4.2.1 Quantitative Verification
(a) For CDEs, we simulate time series data using the predicted functions and assess their numerical fitness using the averaged coefficient of determination: $R^{2}=1/D\sum_{j}^{D}(1-\frac{\sum_{t=1}^{T}\|x^{t}_{j}-\hat{x}^{t}_{j}\|^{2}}{\sum_{t=1}^{T}\|x^{t}_{j}-\bar{\mathbf{x_{j}}}\|^{2}}),$ where $x^{t}_{j}$ denotes the ground truth at time $t$ for time series $j$, $\hat{x}^{t}_{j}$ the model prediction, and $\bar{\mathbf{x}_{j}}$ the empirical mean of the observed data. We also evaluate expression complexity based on symbolic structure, defined as the number of operations. (b) For Boolean networks, we simulate transitions from the predicted Boolean networks and evaluate predictive performance using the macro-averaged F1 score over $T$ transitions: $\mathrm{F1}=\frac{1}{D*T}\sum_{j=1}^{T}\sum_{i=1}^{D}\frac{2\cdot\mathrm{TP}^{j}_{i}}{2\cdot\mathrm{TP}^{j}_{i}+\mathrm{FP}^{j}_{i}+\mathrm{FN}^{j}_{i}},$ where $\mathrm{TP}^{j}_{i}$, $\mathrm{FP}^{j}_{i}$, and $\mathrm{FN}^{j}_{i}$ are the true positives, false positives, and false negatives for node $i$ at transition $j$, respectively. Expression complexity is computed analogously to CDEs. (c) For structural causal models, where the true data-generating process is typically unknown, we adopt a Conditional Independence score (CI-score) as a proxy for structural fidelity. This metric captures the strength of dependence between each child node and its parents, conditioned on the remaining parents. Let the parent set of $X^{i}$ defined as $P_{i}=\{(\mathbf{x}_{j},\ell)|j\in D,\ell\in\ell_{\max}\}$ where $\ell_{\max}$ is the maximum time lag. For each directed edge $(\mathbf{x}_{j},\ell)\to i$ in the edge set $E$, we compute the partial correlation: $\scriptstyle r_{j\to i}=\operatorname{Corr}\!\left(X^{t-\ell}_{j},\;X^{t}_{i}\;\middle|\;\left\{x^{t-\ell^{\prime}}_{k}:(\mathbf{x}_{k},\ell^{\prime})\in P_{i},\;k\neq j\right\}\right).$ The CI-score for a candidate graph $\mathcal{G}$ is then given by: $\mathrm{CI\text{-}score}(\mathcal{G})=\frac{1}{|E|}\sum_{(\mathbf{x}_{j},\ell)\to i\in E}\left|r_{j\to i}\right|.$ A lower CI-score suggests stronger conditional independence and thus better structural plausibility.
4.2.2 Qualitative Verification
To assess the symbolic quality of the generated expressions, we employ a rubric-based scoring system ranging from 1 (poor) to 5 (excellent). Each qualitative score is assigned by an LLM acting as a judge Hao et al. (2024), using a standardized evaluation rubric. The rubric consists of four core criteria: (a) Context alignment: Alignment with provided time series data and contextual descriptions; (b) Scientific plausibility: Alignment with possible physical laws or domain-specific constraints; (c) Conciseness and clarity: The readability and succinctness of the reasoning path; (d) Logical coherence: The consistency and step-by-step soundness of the derivation process.
4.3 Context Manager
The Context Manager is responsible for delivering relevant contextual information to the Proposal Generator during iterative refinement. Based on the chosen refinement strategy, the Context Manager selects and provides appropriate context for the next round of proposal generation. Throughout the process, the Context Manager maintains a history pool $\mathcal{H}$, which stores previously generated expressions, reasoning paths, and their associated verification scores in a structured DataFrame, while removing duplicates. In this study, we adopt a naive ranking-based strategy by only selecting the top-$k$ highest-scoring candidates for further refinement.
5 Experiment
5.1 Experimental Setup
Baseline Models. For CDEs, we include the following baselines: PySR Cranmer (2024), a GP-based method; ProGED Omejc et al. (2024), a probabilistic grammar-based approach; and ODEformer d’Ascoli et al. (2023), a pretrained transformer-based model. For BNs, we evaluate LogicGep Zhang et al. (2024), a GP-based method tailored for Boolean Network inference. For SCMs, we adopt several time-series causal discovery methods, including PCMCI Runge et al. (2019), LPCMCI Gerhardus and Runge (2020), and j-PCMCI+ Günther et al. (2023).
Evaluation Metrics. SymbolBench evaluates models on both in-distribution (ID) and out-of-distribution (OOD) data by applying new initial conditions. For CDEs, we report symbolic regression score: $SR^{2}=\frac{1}{N}\sum_{i=1}^{N}R_{i}^{2}\cdot\mathbb{I}(R_{i}^{2}>0)$, and accuracy for $R^{2}>0.9$: $ACC_{0.9}$. For BNs, we assess numerical similarity via precision, recall, F1, and bookmaker informedness: recall+specificity1recall+specificity-1. For SCMs, structural accuracy is measured using classification metrics and Structural Hamming Distance (SHD). Symbolic similarity (for CDEs and BNs) is evaluated via expression tree edit distance, and expression complexity via sympy’s count_ops Meurer et al. (2017). For BNs and SCMs, we also report accuracy for samples with F1 above a threshold: $ACC_{thesh}$.
5.2 Experimental Results
Across all three tasks, we select LLMs with various sizes and architectures, including Qwen2.5-14B Team (2024), Llama-3.2B Dubey et al. (2024), Mathstral-7B Jiang et al. (2023), GPT-4o-mini Achiam et al. (2023), and ChatTS-14B Xie et al. (2024). For GPT-4o-mini and ChatTS-14B, we use inputs with visual and temporal modalities. We present the results in Table 2,  3, and  4.
Obs. 1: LLMs demonstrate superior capability compared to baselines on CDEs and SCMs datasets, while failing to compete against baselines in Boolean network inference. For Boolean network inference, while LLMs consistently achieve positive bookmaker informedness scores, indicating performance better than random guessing, the genetic programming-based model, LogicGep, significantly outperforms all evaluated LLMs across nearly all evaluation metrics in both ID and OOD scenarios. This could be explained by both the symbolic and numerical fitting process. (a) Symbolic fitting Unlike CDEs, BN inference does not involve coefficient optimization. Compared with CDEs, even if the symbolic structure is not precisely correct, adjusting coefficients can still yield good numerical accuracy. Compared with SCMs, SCM inference focuses only on discovering causal relationships, which is a less stringent goal than recovering the full dynamics as in BN inference. (b) Numerical fitting: CDEs and SCMs are inferred from continuous time series, which form a consistent chains of state transitions, whereas BN inference uses state transitions with various initial conditions, forming graphs of state transitions, as shown in Appendix B and  H.1, which can be hard to identify trends and summarize patterns.
Obs. 2: LLMs’ performance degrades with problem difficulty. Across all models, performance consistently declines as the dimensionality of the system increases. While conventional methods such as PySR remain competitive in low-dimensional systems (e.g., dim = 1), baseline models tend to exhibit a steeper performance drop compared to LLMs. Similarly, there is a notable drop in accuracy when models are evaluated on OOD data, underscoring the increased complexity and generalization challenges posed by these scenarios.
Obs. 3: Chain-of-thought prompting does not consistently improve performance. Further introducing CoT prompting does not lead to consistent gains, especially on the CDEs dataset. An exception is ChatTS-14B, likely due to its specialized time series reasoning capabilities. While increasing test-time compute has been shown to improve outcomes in various tasks Snell et al. (2024), and CoT can contribute to this effect, similar limitations have been observed in more complex tasks such as those in SciBench Wang et al. (2023).
Obs. 4: Providing problem contexts improves LLM performance. Across all three tasks, compared to the Naive prompt, LLMs demonstrate the ability to leverage the provided context, resulting in higher numerical performance. In this setting, context functions as a form of conditioning, helping to constrain the solution space and guide the model toward more accurate and relevant inferences with a faster convergence rate, as shown in Appendix G.
Table 2: Symbolic regression performance for CDEs across 4 dimensions. We use percentage for $SR^{2}$ and $ACC_{0.9}$. Yellow, Orange, and Cyan mark the first, second, and third place, respectively.
Dim Metric Baselines Qwen2.5 Llama3.2 Mathstral 4o-text 4o-image ChatTS
Pro. PySR ODE. Naïve Base Ctx CoT Ctx CoT Ctx CoT Ctx CoT Ctx CoT Ctx CoT
Dim=1
Complexity $\downarrow$ $5.06$ $2.91$ $4.68$ $4.17$ $5.14$ $5.34$ $3.91$ $1.89$ $1.97$ $2.63$ $2.69$ $2.31$ $2.36$ $2.37$ $2.09$ $3.29$ $3.00$
Symbolic Prox. $\downarrow$ $4.86$ $3.54$ $5.18$ $5.11$ $5.37$ $5.57$ $4.89$ $4.03$ $4.14$ $3.86$ $4.09$ $3.89$ $3.79$ $3.97$ $4.06$ $4.83$ $4.37$
ID $SR^{2}$ $95.50$ $96.90$ $83.40$ $93.30$ $97.20$ $99.00$ $97.40$ $95.10$ $94.90$ $96.20$ $96.20$ $92.80$ $95.40$ $95.70$ $95.30$ $95.10$ $97.00$
$ACC_{0.9}$ $91.40$ $97.10$ $73.50$ $91.40$ $97.10$ $100.00$ $97.10$ $91.40$ $91.40$ $97.10$ $97.10$ $94.30$ $97.00$ $97.10$ $97.10$ $91.40$ $97.10$
OOD $SR^{2}$ $74.10$ $88.20$ $54.10$ $60.90$ $74.00$ $61.80$ $61.80$ $69.90$ $64.60$ $71.00$ $66.90$ $70.80$ $66.20$ $71.50$ $62.90$ $63.80$ $70.30$
$ACC_{0.9}$ $65.70$ $85.70$ $47.10$ $45.70$ $63.60$ $52.90$ $51.40$ $51.40$ $48.60$ $62.90$ $60.00$ $54.30$ $51.50$ $54.30$ $45.70$ $51.40$ $54.30$
Dim=2
Complexity $\downarrow$ $5.71$ $7.00$ $9.05$ $8.76$ $9.75$ $10.00$ $9.07$ $6.93$ $7.07$ $7.93$ $8.35$ $6.91$ $6.64$ $6.33$ $6.33$ $9.88$ $9.63$
Symbolic Prox. $\downarrow$ $10.90$ $8.21$ $9.98$ $11.20$ $11.60$ $11.40$ $11.10$ $10.00$ $9.86$ $10.30$ $10.40$ $10.40$ $9.98$ $10.10$ $10.00$ $11.70$ $11.40$
ID $SR^{2}$ $61.20$ $80.40$ $72.50$ $82.40$ $87.90$ $87.40$ $88.00$ $82.20$ $80.90$ $88.60$ $88.40$ $78.50$ $77.40$ $78.40$ $82.40$ $84.80$ $90.90$
$ACC_{0.9}$ $42.90$ $76.90$ $47.60$ $71.10$ $77.30$ $80.00$ $76.70$ $65.10$ $65.10$ $79.10$ $79.10$ $62.20$ $59.50$ $62.80$ $62.80$ $76.70$ $83.70$
OOD $SR^{2}$ $18.80$ $54.00$ $34.60$ $44.90$ $52.90$ $45.30$ $49.70$ $40.60$ $47.30$ $46.10$ $46.80$ $42.70$ $40.20$ $38.10$ $45.60$ $49.50$ $44.70$
$ACC_{0.9}$ $12.50$ $42.10$ $18.60$ $31.00$ $44.40$ $34.10$ $35.90$ $32.60$ $34.90$ $34.90$ $33.30$ $29.50$ $22.50$ $23.80$ $29.30$ $43.60$ $33.30$
Dim=3
Complexity $\downarrow$ $7.87$ $8.82$ $13.90$ $16.60$ $17.40$ $18.30$ $15.50$ $15.70$ $17.30$ $16.30$ $15.40$ $12.60$ $12.30$ $12.00$ $11.50$ $17.00$ $17.20$
Symbolic Prox. $\downarrow$ $30.30$ $25.70$ $30.50$ $29.80$ $30.10$ $29.70$ $28.60$ $28.80$ $30.00$ $29.50$ $29.00$ $29.50$ $27.40$ $29.20$ $29.20$ $29.40$ $29.60$
ID $SR^{2}$ $20.90$ $67.10$ $36.30$ $71.30$ $74.30$ $75.30$ $70.90$ $62.10$ $57.40$ $65.20$ $64.60$ $70.00$ $63.70$ $65.30$ $66.90$ $70.80$ $68.40$
$ACC_{0.9}$ $10.90$ $60.50$ $18.40$ $61.20$ $63.30$ $67.30$ $53.10$ $49.00$ $43.50$ $57.10$ $57.10$ $56.30$ $50.00$ $49.00$ $53.10$ $59.20$ $57.10$
OOD $SR^{2}$ $8.60$ $49.90$ $30.00$ $47.20$ $43.20$ $44.30$ $40.90$ $43.10$ $35.90$ $38.70$ $45.00$ $44.70$ $39.40$ $39.20$ $37.30$ $40.20$ $42.40$
$ACC_{0.9}$ $4.50$ $36.10$ $15.20$ $39.50$ $38.60$ $36.60$ $33.30$ $34.00$ $28.60$ $33.30$ $40.90$ $38.30$ $31.70$ $31.10$ $29.20$ $32.60$ $33.30$
Dim=4
Complexity $\downarrow$ $9.70$ $10.60$ $14.40$ $22.90$ $23.20$ $23.10$ $18.50$ $20.60$ $19.60$ $17.90$ $18.70$ $14.50$ $13.90$ $13.90$ $13.60$ $22.60$ $22.10$
Symbolic Prox. $\downarrow$ $32.70$ $30.30$ $36.40$ $35.20$ $35.40$ $34.10$ $32.90$ $34.70$ $32.70$ $32.50$ $32.40$ $34.80$ $32.90$ $32.20$ $31.90$ $38.00$ $36.30$
ID $SR^{2}$ $11.50$ $74.60$ $34.80$ $92.70$ $93.60$ $93.30$ $91.00$ $49.40$ $61.10$ $78.00$ $78.70$ $88.60$ $86.90$ $86.10$ $89.10$ $78.70$ $80.00$
$ACC_{0.9}$ $8.70$ $65.20$ $29.60$ $87.90$ $94.10$ $93.90$ $89.30$ $40.70$ $52.00$ $74.10$ $75.00$ $78.80$ $70.00$ $71.40$ $85.70$ $74.10$ $80.80$
OOD $SR^{2}$ $7.00$ $23.80$ $11.40$ $29.20$ $19.10$ $19.40$ $34.60$ $28.30$ $21.30$ $30.70$ $36.90$ $26.90$ $27.60$ $30.40$ $30.50$ $25.00$ $24.50$
$ACC_{0.9}$ $0.00$ $21.70$ $7.10$ $19.40$ $15.60$ $12.50$ $33.30$ $12.00$ $13.60$ $24.00$ $30.80$ $19.40$ $20.70$ $15.40$ $21.40$ $21.70$ $17.40$
Table 3: Comparison of Boolean network inference across ID and OOD settings.
Model Setting ID OOD Symb. Prox. $\downarrow$ Comp. $\downarrow$
Prec. Rec. Acc. B.I. $\mathbf{ACC_{0.5}}$ $\mathbf{ACC_{0.7}}$ $\mathbf{ACC_{0.8}}$ Prec. Rec. Acc. B.I. $\mathbf{ACC_{0.5}}$ $\mathbf{ACC_{0.7}}$ $\mathbf{ACC_{0.8}}$
LogicGep $\sim$ 93.6 92.7 95.2 88.7 98.5 98.5 96.9 84.7 86.5 89.5 76.5 98.5 86.2 76.9 12.39 12.39
Qwen2.5-14B Naïve 58.7 71.7 66.5 29.5 86.2 33.8 6.2 56.4 71.3 64.5 26.6 80.0 24.6 7.7 12.87 14.84
Base 54.7 73.3 63.5 27.3 80.0 20.0 3.1 53.8 72.8 62.5 25.9 83.1 16.9 3.1 12.33 16.73
Context 57.8 77.2 67.1 29.1 87.7 38.5 9.2 56.1 77.2 65.5 27.0 87.7 30.8 10.8 12.39 16.86
CoT 58.3 73.9 65.4 30.3 92.3 27.7 4.6 56.2 73.9 63.2 27.9 84.6 24.6 3.1 11.96 14.79
Llama3.2-3B Context 51.6 74.8 62.4 24.5 60.0 18.5 1.5 47.4 71.1 58.2 16.9 52.3 9.2 1.5 14.12 17.67
CoT 52.0 73.1 62.3 23.0 61.5 20.0 1.5 49.9 67.8 58.3 15.7 56.9 13.8 1.5 14.30 19.59
Mathstral-7B Context 49.5 73.3 60.3 19.2 50.8 15.4 1.5 48.0 72.2 58.0 16.4 53.8 10.8 1.5 12.74 20.14
CoT 51.3 74.4 62.6 23.9 61.5 21.5 3.1 48.3 71.7 58.3 16.9 60.0 16.9 3.1 12.16 20.48
GPT-4o-mini Context 51.6 58.9 62.6 21.2 47.7 15.4 0.0 51.3 58.5 61.4 20.0 53.8 13.8 0.0 13.24 30.07
CoT 52.2 58.9 63.5 23.5 53.8 4.6 0.0 51.7 60.6 62.2 23.0 53.8 7.7 0.0 11.20 32.42
Table 4: Performance comparison of causal discovery methods and LLM-based approaches.
[…]
PCMCI $\sim$ 52.7 52.2 63.7 46.1 55.3 23.7 7.9 95.28 96.76
LPCMCI $\sim$ 52.0 68.6 45.9 29.1 56.3 18.4 5.3 25.35 14.54
j-PCMCI+ $\sim$ 46.2 58.1 43.3 39.1 45.3 13.2 7.4 49.36 38.83
Qwen2.5-14B Naïve 49.7 59.4 45.0 40.6 43.7 17.9 8.4 35.54 18.93
Base 50.5 61.4 45.0 38.6 47.4 18.4 8.9 34.58 18.10
Context 53.4 63.0 48.3 37.0 53.7 19.5 7.9 31.69 18.61
CoT 51.3 61.9 46.3 38.1 52.6 20.0 8.4 39.36 21.76
Llama3.2-3B Context 51.3 56.1 49.0 43.9 37.4 13.7 8.9 41.89 25.43
CoT 51.8 56.4 49.5 43.6 48.4 22.6 11.6 44.40 25.75
Mathstral-7B Context 47.2 54.8 43.1 45.2 40.5 15.8 7.9 37.70 20.48
CoT 49.8 58.0 45.4 42.0 44.7 14.2 5.8 34.30 19.38
GPT-4o-mini Context 36.9 59.9 27.6 40.1 24.7 5.3 1.6 44.29 13.36
CoT 37.1 57.5 28.5 42.5 24.2 3.2 2.6 38.90 13.08
ChatTS-14B Context 54.1 72.3 46.2 27.7 58.9 22.6 11.6 32.48 15.07
CoT 54.4 72.3 46.6 27.7 61.1 25.3 10.5 31.10 14.70
Table 5: Correlations between complexity and OOD $ACC_{0.9}$ across four dimensions with and without context in symbolic regression task.
Condition Dim 1 Dim 2 Dim 3 Dim 4
w/o context 0.672-0.672 0.384-0.384 0.425-0.425 $0.584$
w/ context $0.097$ $0.728$ $0.165$ 0.174-0.174
5.3 Further Analysis
In this section, we explore the adaptability of our framework with genetic programming methods and the power of test-time compute. We formulate several key questions and takeaways as follows:
Q1: Do LLMs Employ Correct Symbolic Reasoning Paths? While previous research Hao et al. (2024) has demonstrated that yielding correct answers does not necessarily yield correct reasoning paths, we find that general LLMs like Qwen2.5-14B with pure textual input are able to perform a certain level of reasoning (through CoT prompting) over the time series and the given context. As shown in Table H.2, the LLM is not only able to consider the meaning of each variable, but also the historical candidates.
Q2: How should test-time compute be structured for consistent improvement? While scaling test-time compute can enhance LLM performance Snell et al. (2024), our findings indicate that the structure of this computation is critical for achieving consistent improvement. Simply increasing compute via naive CoT reasoning does not yield reliable gains, as shown in Obs. 3. This is because the depth of naive CoT reasoning often remains shallow and lacks key mechanisms such as reflection, verification, and backtracing that are present in more advanced RLM Besta et al. (2025) with Long CoT Chen et al. (2025), which involves more cognitive behaviors such as verification, reflection, and backtracing, etc. We use illustrative examples to demonstrate this in Appendix H.3.
Thus, to see consistent gains, test-time compute should be structured to facilitate a Long CoT. We explore two approaches: First, we analyze the effect of a Long CoT, generated by RLMs, within each reasoning epoch. This reveals that dedicating more compute to a longer, more detailed chain-of-thought leads to moderate but consistent improvements in the final prediction scores, as shown in Figure 3. Second, we view the entire iterative refinement process as a form of Long CoT. In this approach, the model verifies previous outputs and generates improved answers based on earlier, potentially flawed, solutions. As shown in Figure 2, the verification scores consistently improve as more refinement steps are added. From both approaches, we observe that the complexity of the best-fitting solutions increases with the amount of computation, echoing the human-like process of progressively constructing more sophisticated answers.
Figure 2: Evaluation scores improve with more iterations and test-time compute, as discussed in Q2.
(a) Comparison of increasing test-time compute on CDEs.
(b) Comparison of w and w/o thinking on BNs and SCMs.
Figure 3: Comparison of performance with and without thinking.
Table 6: Hybrid method on CDEs. We use the same GPT-4o-mini as judge and predictor.
Method $SR_{ID}^{2}$ $SR_{OOD}^{2}$ Sym. Prox. $\downarrow$ Comp. $\downarrow$
GPLearn 27.6 14.9 6.083 1.333
GPT-4o-mini 39.3 24.1 7.438 7.406
GPLearn + LLM-as-Judge 31.7 16.3 6.550 1.200
LLM-as-Predictor + GPLearn 89.5 69.3 5.045 2.682
Table 7: Hybrid method on BNs. We use the same GPT-4o-mini as judge and predictor.
Method Setting Metrics
Prec. Recall F1 Acc BM Comp. $\downarrow$
LogicGep ID 88.0 88.6 88.0 91.7 80.9 0.773
OOD 78.9 83.0 80.0 85.6 69.2 0.773
GPT-4o-mini ID 52.8 65.6 57.8 61.2 23.6 2.422
OOD 49.9 62.3 54.5 58.3 17.3 2.422
LogicGep + Judge ID 89.3 92.8 90.9 93.2 85.6 1.012
OOD 78.6 85.2 80.4 85.8 71.0 1.012
Predictor + LogicGep ID 72.2 84.7 77.5 80.6 61.0 0.993
OOD 63.1 77.2 67.9 73.2 46.9 0.993
Q3: Does the generalizability of prediction correlate with structural complexity? When faced with multiple symbolic structures that can fit the given time series, the conventional wisdom suggests that the simplest one is the most likely to be generalizable. Existing LLM-based approaches for symbolic regression Shojaee et al. (2024); Li et al. (2024); Merler et al. (2024); Grayeli et al. (2024); Wang et al. (2025), rooted in the principle of Occam’s Razor, advocate for choosing the expression with the least complexity. However, our experimental results challenge the assumption that simpler is always better. Our analysis shows that expressions with higher complexity can also achieve improved symbolic proximity and OOD performance in Table 2. This is because, in complex scientific domains such as biology, physics, and healthcare Xie et al. (2025), the ground-truth expression may inherently be more structurally complex while being numerically simpler in the ID setting. Imposing a strong simplicity criterion can hinder the discovery or selection of the more generalizable context-aligned candidates. As detailed in Appendix E and Table 5, our analysis suggests that: Candidate ranking and selection should not solely rely on complexity and may benefit from considering contextual information.
Q4: Can LLMs and genetic programming (GP) be combined for enhanced performance? While LLMs show superior performance on CDE and SCM inference, GP holds value for its cost, control over complexity, and stronger performance in BN inference. To capitalize on the advantages of both, we explored two primary methods for creating a high-performance hybrid system, as detailed in Appendix F. The key is to use the contextual understanding of LLMs to “inject context and knowledge” into the GP workflow, boosting its performance. This can be achieved in two main ways: (a) LLM-Guided Initialization: Use an "LLM-as-Predictor" to generate a high-quality initial population of candidate expressions for the GP to evolve. This provides the GP with a context-aware and promising starting point. (b) LLM-Guided Evaluation: Employ an “LLM-as-Judge” to provide context-enhanced evaluation scores during the GP’s fitness assessment. This helps guide the evolutionary search toward solutions that are not only numerically accurate but also scientifically plausible. As shown in Tables 6 and 7, combining LLMs with GP in these roles demonstrably improves performance.
6 Conclusion and Outlook
We introduce SymbolBench, a real-world benchmark for symbolic structure discovery, and a Unified Symbolic Reasoning Framework that enables LLMs (optionally with GP) to generate and judge hypotheses across tasks. Experiments show: (i) LLMs beat baselines on multivariate symbolic regression and causal discovery but lag on Boolean network inference; (ii) more test-time compute yields only modest gains; (iii) contextual grounding boosts accuracy and generalizability; and (iv) LLM–GP hybrids further improve performance.
Future Opportunities. Based on our observations and takeaways, we summarize some potential future directions as follows: (i) task-specific scaling of test-time compute and reasoning depth; (ii) richer, knowledge-heavy context to guide hypotheses; (iii) context-aware criteria beyond syntactic simplicity (e.g., plausibility, robustness, causal faithfulness); and (iv) broader symbolic targets plus verifiable reasoning traces for interpretability.
7 Limitations
This work has several limitations stemming from computational constraints. First, we did not evaluate larger open-source models (e.g., DeepSeek-R1). Second, we capped the budget at 100 generation epochs and a maximum of 20 retries per epoch for each run; samples without candidates that could reach the tolerance will be rerun once again to mitigate uncertainty. Finally, we examined only one LLM and one GP in the hybrid architecture, and did not analyze the effects of reasoning strategies or model size.
Appendix A Related Work
symbolic structure Discovery. Discovering symbolic laws from time series data is a central objective in many scientific discovery tasks. One prominent approach is Symbolic Regression, which seeks closed-form expressions that accurately model the observed data. Classical methods such as Genetic Programming (GP) Mundhenk et al. (2021) are powerful but computationally intensive and often sensitive to the choice of operators and fitness functions. Alternatively, sparse optimization techniques like SINDy Brunton et al. (2016) and PySR Cranmer (2024) aim to identify parsimonious models by leveraging sparsity in the function space. More recently, deep learning-based models have been introduced to enhance efficiency and scalability. Methods such as ODEformer d’Ascoli et al. (2023) and TPSR Shojaee et al. (2023) reformulate symbolic discovery as a sequence-to-sequence translation task, mapping time series data to symbolic equations. These approaches generate high-quality expressions and offer improved computational efficiency. However, they typically require large-scale pretraining and often lack the capability for iterative refinement and adaptation across diverse scientific domains. Beyond symbolic regression, related tasks such as Boolean Network InferenceZhang et al. (2024) and Causal DiscoveryHasan et al. (2023) also aim to extract symbolic structures from time series data. These methods seek to uncover underlying logical or causal relationships, further emphasizing the broader interest in interpretable, symbolic representations of dynamical systems.
LLM Symbolic Reasoning. Due to the strong in-context learning and reasoning ability of Large language models (LLMs) that allow them to adapt to various tasks, they are able to perform logical inferenceAhn et al. (2024); Wang and Chen (2023), and temporal symbolic reasoning Fang et al. (2024). Despite the strong ability of LLMs, current research for SR tasks only applies them as proposal generators without a reasoning process. Though the generated functions may achieve a high fitting score, the reasoning process remains unknown to us and may fail to align with the context, with little real-world meaning. In addition, the current research focus has also skewed toward algebraic equations, with little work on logic rules or causal relations. Our SymbolBench, compared to previous benchmarks as shown in Table 1, addresses this by systematically evaluating both LLMs and MLLMs on SR for time series data, covering varied symbolic forms and emphasizing the transparency of the reasoning process. In addition, compared to prior benchmarks, we provide a more comprehensive evaluation including In-distribution and Out-of-distribution settings over coupled ODEs, logical expressions, and structured causal models.
Appendix B SymbolBench Dataset
SymbolBench uses three different datasets to evaluate the ability of LLMs/MLLMs to uncover the symbolic laws from time series data. We include variable description, domain name, and the time series trajectory as additional context. Examples of the dataset are shown in Table 8.
Table 8: Summary of datasets, ground truths, and variable context with trajectory images.
Dataset Example GroundTruth Variable Description Domain Trajectory
Coupled Differential Equation dx1/dt = (c0 + c1 / (1 + c2 * x2)) - c3 * x1 x1: ng_ml Endocrine [Uncaptioned image]
dx2/dt = (c8 + c9 * x3) - c10 * x2 x2: microg_dl
dx3/dt = (c4 + c5 * x1) / (1 + c6 * x2) - c7 * x3 x3: pg_ml
Boolean Network x1 = ( NOT ( x3 OR x5 ) OR NOT ( x5 OR x3 ) ) x1: v_Coup_fti Cortical Area Development [Uncaptioned image]
x2 = ( x1 AND NOT ( ( x3 OR x5 ) OR x4 ) ) x2: v_Emx2
x3 = ( ( x3 AND x5 ) AND NOT x2 ) x3: v_Fgf8
x4 = ( x5 AND NOT ( x2 OR x1 ) ) x4: v_Pax6
x5 = ( x3 AND NOT x2 ) x5: v_Sp8
Structured Causal Model $\sim$ x0: membrane (millivolt) Calcium Dynamics [Uncaptioned image]
$\sim$ x1: rapidly_activating_K_current_n_gate (dimensionless)
$\sim$ x2: Ca_i in component ionic_concentrations (micromolar)
$\sim$ x3: None
In addition, we provide a more detailed illustration of the verification process of the three tasks as follows:
  1. (a)
    CDEs: We use an LLM to generate the skeleton of the continuous-time dynamical system, replacing each unknown coefficient with a placeholder "c". An optimizer then fits these coefficients to the training time series. To avoid the high cost of repeatedly solving and differentiating through a full ODE solver, we adopt the finite–difference approximation strategy from ODEFormer d’Ascoli et al. (2023). Once the functional form $\hat{f}(\mathbf{x};\boldsymbol{\phi})$ is obtained, where $\phi$ is the fitted coefficients, we generate the final numerical solution using SciPy’s scipy.integrate.solve_ivp, and compare the trajectory to the ground truth:
    $\texttt{solve\_ivp}\bigl{(}\hat{f}(\mathbf{x};\boldsymbol{\phi}),\;\mathbf{x}(t_{0}),\;t_{0},\ldots,t_{n},\texttt{method=LSODA}\bigr{)}.$ (1)
    The distribution of sample dims is shown in Figure 4(a).
  2. (b)
    BNs: For Boolean Networks, the LLM directly outputs a set of logical update rules (e.g. $x_{i}(t+1)=x_{j}(t)\,\land\,\lnot x_{k}(t)$). Since there are no continuous parameters to fit, we simply simulate the network from the known initial state $x(t_{0})$ and compute the F1 score over all bits and time steps to assess agreement with the true dynamics. The distribution of sample dims is shown in Figure 4(b).
  3. (c)
    SCMs: For Structured Causal Models, the LLM predicts the possible causal relations among all variables, forming a directed graph. Since the predicted SCMs can not directly produce numerical solutions, to quantify how well the predicted SCM explains the data, we compute the sample partial correlation between $x_{i}$ and each candidate parent in $\mathrm{pa}(x_{i})$ (conditioning on the parents), following the protocol of Runge et al. Runge et al. (2019). The final score for each node is the mean of its absolute partial correlations, and we average over all nodes to obtain the overall SCM score. The distribution of sample dims is shown in Figure 4(c).
Textual Context. For regular LLMs that only accept textual inputs, all inputs to LLMs are formatted as structured textual prompts. Time series data are serialized into strings, supplemented with contextual metadata such as domain information, variable meanings, and prior scored expressions. This enables language models to reason over both data patterns and contextual priors.
Visual and Temporal Context. Multimodal LLMs have shown promise in handling visual and temporal data. While visual inputs are traditionally used in vision-language tasks Radford et al. (2021), recent research has extended MLLMs to time series domains (e.g., forecasting via visual encodings Li et al. (2023)). Recent models like ChatTS Xie et al. (2024) and TempoGPT Zhang et al. (2025) enable joint reasoning over temporal and textual modalities. In SymbolBench, we explore the use of MLLMs to incorporate both visual time series plots and encoded temporal embeddings.
(a) Dimension distribution of Coupled Differential Equations used in SymbolBench.
(b) Dimension distribution of Boolean Networks used in SymbolBench.
(c) Dimension distribution of Structured Causal Models used in SymbolBench.
Appendix C LLMs for Benchmarking
We evaluate six representative LLMs, chosen for their diversity in size, training data, and specialization:
  1. (a)
    Qwen Series Team (2024, 2025): Qwen2.5-14B Team (2024) is a 14.7-billion-parameter causal Transformer (13.1 B non-embedding) built on RoPE, SwiGLU, RMSNorm, and QKV-bias that extends context support to 128 K tokens (with generation up to 8 K) and delivers significantly richer knowledge, advanced coding, and mathematical reasoning (via domain-expert submodels), robust instruction following, long-form text, and structured output (e.g., tables, JSON), and out-of-the-box multilingual fluency across 29+ languages. Following Qwen2.5-14B, Qwen3 Team (2025) series is the newest generation in the Qwen family, combining dense and Mixture-of-Experts architectures to deliver seamless mode-switching—“thinking” for deep logical reasoning, math, and coding, and “non-thinking” for fast, general dialogue.
  2. (b)
    Llama 3.2-3B Dubey et al. (2024): Llama 3.2 is a family of multilingual 1 B and 3 B–parameter pretrained and instruction-tuned text-in/text-out models, with its instruction-tuned versions specially optimized for dialogue, agentic retrieval, and summarization across dozens of languages—consistently outperforming many open-source and proprietary chat models on standard industry benchmarks.
  3. (c)
    Mathstral-7B Jiang et al. (2023): Mathstral is a 7 billion-parameter LLM released by Mistral AI as a tribute to Archimedes’ 2311th anniversary, built on Mistral 7B with a 32K token context window and fine-tuned for advanced multi-step mathematical and scientific reasoning. Developed in collaboration with Project Numina, it achieves state-of-the-art performance for its size.
  4. (d)
    ChatTS-14B Xie et al. (2024): ChatTS-14B is a multimodal LLM explicitly designed around time series as its core modality, offering native support for multivariate sequences of varying lengths and dimensions, preserving raw numerical fidelity for precise statistical queries, and enabling interactive, conversational exploration and reasoning over time-series data—while also integrating seamlessly into existing LLM workflows (including vLLM) with provided code, datasets, and models.
  5. (e)
    GPT-4o-mini Achiam et al. (2023): GPT-4o-mini is a compact multimodal reasoning model released by OpenAI in July 2024, delivering GPT-4–level performance while costing over 60% less than GPT-3.5 Turbo; it supports text and vision inputs, advanced function calling, and extended long-context understanding.
  6. (f)
    o4-mini Achiam et al. (2023): o4-mini is OpenAI’s latest release of a reasoning-focused GPT variant that replaces o3-mini, offering both text and image processing, “whiteboard” chain-of-thought reasoning, seamless tool integration, and a high-accuracy paid-tier option—all accessible via ChatGPT and the Completions API for domain-critical decision-making tasks.
Appendix D Baseline Implementation
  1. (a)
    CDEs: We follow ODEFormer’s baseline implementation and hyperparameter tuning d’Ascoli et al. (2023), using PySR with finite-difference approximations for skeleton search, and default greedy top-$k$ generation for ODEFormer.
  2. (b)
    BNs: We reimplement LogicGep’s Boolean-network inference using the same Geppy genetic-programming framework Gao et al. (2022), but omit the continuous-to-binary discretization and MLP-based constraint stages, since our training traces are already binary.
  3. (c)
    SCMs: All SCM baselines are based on Tigramite Runge et al. (2022), with the maximum time-lag set to 1 for fair comparison.
Appendix E Analysis on Table 2
Prior studies Shojaee et al. (2024); Merler et al. (2024) often use complexity as the sole standard for final selections of candidate predictions. However, we show that the generalization, represented by the performance during holdout evaluation (OOD), has a poor correlation with expression complexity. (a) As shown in Table 9, without introducing context, complexity may have a moderate correlation with $ACC_{0}.9$ when the dimension is small and not challenging. However, given samples with higher dimensions, the correlation became positive, meaning higher complexity can also give better generalization. (b) When context is introduced, besides the overall improved performance is observed as illustrated in Obs. 4, the correlation also turned positive from Dim=1 to Dim=3 and remained low for Dim=4. Such an obvious change potentially suggests that context is a more effective criterion for candidate ranking and selection during both iterative refinement and final evaluation.
Table 9: Correlations between complexity and OOD $ACC_{0.9}$ across four dimensions with and without context
Condition Dim 1 Dim 2 Dim 3 Dim 4
w/o context 0.672-0.672 0.384-0.384 0.425-0.425 $0.584$
w/ context $0.097$ $0.728$ $0.165$ 0.174-0.174
Appendix F Hybrid Method
We adopt two frameworks for hybrid approach by separately let make GP and LLMs play different roles. (a) As shown in Figure 5(a), in addition to the quantitative evaluation using MSE, an additional qualitative score produced by LLM is incorporated in the evolution loop; (b) As shown in Figure 5(b), in addition to the original closed loop (blue line), we provide an extended path (red line) that utilize GP to expand the history pool. From a different perspective, the initial population produced by LLMs also improves the generation for GP by providing context-enhanced initial populations.
(a) Hybrid method using Genetic Programming + LLM-as-Judge.
(b) Hybrid method using Genetic Programming + LLM-as-Predictor, where GP helps expand the history pool with the current best expressions as initial population.
Figure 5: Hybrid method
Appendix G Convergence Rate
We further examine the convergence rate under various settings. As shown in Figure 6, introducing both context and reasoning leads to a faster convergence rate on the CDE and BN datasets. In contrast, the SCM dataset exhibits an overall fast convergence rate across all settings, with context having only a marginal improvement.
(a) CDE
(b) BN
(c) SCM
Figure 6: Convergence Rate on CDE, BN, and SCM datasets.
Appendix H Example Outputs
H.1 Example predictions across three tasks
We present examples across three tasks in Table 8.
Figure 7: Example input and output across three tasks using LLMs-as-Predictors. The output is processed through verification.
H.2 Candidates of CoT reasoning output on coupled-differential equations
We show the list of candidates found for fitting the time series generated from the SEIR model in Table 10.
Table 10: Example candidates of a 4-dimensional coupled differential equation during inference.
# expr R2 Complexity reasoning (Qwen2.5-14B) context_alignment scientific_plausibility conciseness&clarity logical_coherence
0 {-c*x_0*x_1 - c*x_0*x_2 + c*x_0,
c*x_0*x_1 + c*x_0*x_2 - c*x_1,
c*x_1 - c*x_2 - c*x_3,
-c*x_0 + c*x_2 + c*x_3}
0.99999 24 This model combines previous models. It introduces
complex interaction between states. Transitions are
governed by rate constant c. Susceptible and
recovered states interact with infected/immune.
4 4 3 4
1 {c*x_0*x_1 - c*x_2 + c*x_3,
c*x_1 + c*x_2 - c*x_3,
-c*x_0 + c*x_2 + c*x_3,
c*x_0 - c*x_1 + c*x_3}
0.99976 21 x_0 and x_1 interact to affect x_2. Other variable
pairs interact linearly. This might reflect a chain
of influence among factors—each influencing the
next in a progression.
4 3 2 3
2 {c*x_0 + c*x_1 - c*x_2,
c*x_1 + c*x_2 - c*x_3,
-c*x_0 + c*x_2 + c*x_3,
c*x_0 - c*x_1 + c*x_3}
0.99975 20 This system maintains cyclic balance among variables.
Represents processes like ecological cycling where
each state influences and is influenced in turn,
sustaining continuous transitions.
2 2 3 2
3 {c*x_0*x_1 - c*x_0*x_2,
c*x_0*x_2 - c*x_1,
c*x_1 + c*x_2 - c*x_3,
-c*x_0 + c*x_3}
0.99825 17 A variant epidemiological model. Introduces a slightly
more intricate structure for transitions. Dynamics
still guided by rate constant c. Greater attention
to susceptible and infectious interplay.
4 4 3 4
4 {c*x_0*x_1 - c*x_2 + c*x_3,
-c*x_0*x_2 + c*x_1 + c*x_3,
c*x_0 - c*x_1*x_3 + c*x_2,
-c*x_0*x_2 + c*x_1 + c*x_3}
0.99292 24 All variables appear in every equation. Introduces
nonlinear terms, enabling feedback loops. May model
biochemical systems with mutual influence and
delayed or amplified effects.
4 3 2 3
5 {c*x_0*x_1 - c*x_2,
c*x_1*x_2 - c*x_3,
-c*x_0*x_1 + c*x_2,
c*x_0 - c*x_1 + c*x_3}
0.96474 17 Models disease spread (SIR). x0: susceptible,
x1: exposed, x2: infected, x3: recovered. Each
equation captures standard transitions. E.g.,
x0*x1 for infections; x2 affects x3.
5 4 4 4
6 {c*x_0 - c*x_1*x_2,
c*x_1*x_2 - c*x_3,
-c*x_0 + c*x_2*x_3,
-c*x_0 + c*x_1 + c*x_3}
0.90896 17 Similar to the SIR model with slight variations.
Includes multiplicative terms between states.
x0 and x1 drive infection; x2 and x3 influence
recovery rates. Captures more nuanced behavior.
4 3 4 4
7 {-c*x_0*x_1 - c*x_0*x_2 + c*x_0,
c*x_0*x_1 - c*x_1,
c*x_0*x_2 - c*x_2,
c*x_1 + c*x_2 - c*x_3}
0.79325 20 Basic epidemiological structure (SIR). Variables:
x0: susceptible, x1: infected, x2: recovered,
x3: immune/dead. Rate constant c influences all
transitions between compartments.
4 4 3 4
8 {-c*x_0*x_1 + c*x_0*x_2 - c*x_2,
c*x_0*x_1 - c*x_1 - c*x_2,
c*x_2 - c*x_3,
-c*x_0*x_2 + c*x_1}
0.33899 20 Represents SEIR model. dx0/dt: loss via infection,
dx1/dt: gain from infection, loss via exposure,
dx2/dt: gain from exposure, loss via recovery,
dx3/dt: gain from recovery, potential loss.
4 4 3 4
H.3 RLM output for one sample with marked reasoning steps
(Step 1: Problem Setup) Okay, let’s see. I need to generate a system of coupled ODEs for the calcium_dynamics domain. The variables are $x_{0},x_{1},x_{2},x_{3}$, each with their descriptions as given. The time series data is provided, which includes $x_{0}$ values and the other variables. (Step 2: Initial Observations) First, the variables $x_{0},x_{1},x_{2},x_{3}$ are each in the range of their description. The domain is calcium dynamics, so the equations need to model some process related to calcium concentration. Looking at the time series for $x_{0}$, which is a negative value, and the other variables being positive. Since calcium concentrations are usually positive, maybe the model is about the flow of calcium through the cell membrane, or the impact of different ions on calcium levels. The time series for $x_{1}$ is all zeros except at some points. The $x_{1}$ values are all 0.3, then increase. Maybe this is a transient phase or a step change. For $x_{2}$ and $x_{3}$, they seem to be varying linearly. (Step 3: First Model Attempt) Looking at the equations, perhaps there are some nonlinear terms. The variables can be combined using trigonometric functions. Let me think about a possible model. If I consider a calcium channel model, the rates of change could involve the concentration of ions. For example, the rate of change of $x_{0}$ might depend on $x_{1},x_{2},$ and $x_{3}$. Similarly, $x_{1}$ might relate to $x_{2}$ and $x_{3}$. (Step 4: Linear Combination Trial) But the problem is to create a model that fits the data. Let’s think of a simple model. Suppose that $x_{0}=c(x_{1}+x_{2}+x_{3})$. Similarly, maybe $x_{1}=c(x_{2}-x_{3})$. For $x_{2}$, maybe it’s a linear function: $x_{2}=c(x_{0}+x_{3})$. Similarly, $x_{3}=c(x_{3}-x_{0})$. Putting it all together, the system would be:
c*x1 + c*x2 + c*x3 | c*x2 - c*x1 | c*x0 + c*x3 | c*x3 - c*x0
(Step 5: Plugging in Sample Data) In the first sample: $x_{0}=-43.0$, $x_{1}=0.03$, etc.
$x_{0}=c(0.03+0.1+0.434)=c(0.564)\Rightarrow c=\frac{-43.0}{0.564}\approx-76.5$
$x_{1}=c(0.1-0.434)=c(-0.334)\Rightarrow c=\frac{0.03}{-0.334}\approx-0.09$
Inconsistent values suggest equation mismatch. (Step 6: Revising the Model) Try again with the same structure:
$c=\frac{x_{0}}{x_{1}+x_{2}+x_{3}}=\frac{-43.0}{0.564}\approx-76.5$ $c=\frac{0.03}{0.1-0.434}=\frac{0.03}{-0.334}\approx-0.09$
Still inconsistent. (Step 7: Trying Differential Form) Suppose:
dx0/dt = c*(x1 + x2 + x3)
dx1/dt = c*(x2 - x3)
dx2/dt = c*(x0 + x3)
dx3/dt = c*(x3 - x0)
(Step 8: Numerical Check) With $x_{0}=-43.0,x_{1}=0.03,x_{2}=0.1,x_{3}=0.434$:
dx0/dt = c*0.564
dx1/dt = c*(-0.334)
dx2/dt = c*0.464
dx3/dt = c*43.434
Still inconsistent c values. (Step 9: Considering Other Forms) Maybe use other terms, e.g., absolute values or trigonometric forms.
Step 10: Trying Absolute Values
dx0/dt = |x1 + x2 + x3|
dx1/dt = |x2 - x3|
dx2/dt = |x1 - x3|
dx3/dt = |x3 - x0|
Still not matching data. (Step 11: Simpler Model Revisited) Try:
dx0/dt = c*x1 + c*x2 + c*x3
dx1/dt = c*x2 - c*x3
dx2/dt = c*x0 + c*x3
dx3/dt = c*x3 - c*x0
(Step 12: Integrating into Sample) With $c=1$:
dx0/dt = 0.564
dx1/dt = -0.334
dx2/dt = -42.566
dx3/dt = 43.434
(Step 13: Final Justification) Suggests calcium dynamics via flow of ions. $x_{0}$ as calcium concentration, $x_{1},x_{2},x_{3}$ as inflow/outflow.
(Step 14: Final Model Form and JSON)
dx0/dt = c*x1 + c*x2 + c*x3
dx1/dt = c*x2 - c*x3
dx2/dt = c*x0 + c*x3
dx3/dt = c*x3 - c*x0
(Step 15: Format Verification) Follow pipe-separated format:
{"eq": "c*x_0*x_1 + c*x_2 - c*x_3 | c*x_1/x_0 + c*x_3 | c*x_2 - c*x_0*x_3 | c*x_3 + c*x_0 - c*x_1",  "dim": 4}
(Final JSON Output)
  "eq": "c*x_0*x_1 + c*x_2 - c*x_3 | c*x_1/x_0 + c*x_3 | c*x_2 - c*x_0*x_3 | c*x_3 + c*x_0 - c*x_1",
  "dim": 4
Appendix I Prompts
Figure 8: Prompt with CoT and Context for CDEs.
Figure 9: Prompt with CoT and Context for BNs.
Figure 10: Prompt with CoT and Context for SCMs.
Figure 11: LLM-as-Judge prompt for CDEs.